# Magnetic-field induced momentum-dependent symmetry breaking in CsV3Sb5 revealed by magneto-ARPES

### v1

@@ III Field response of the Sb 5 band at
Next, we turn our attention to the electron pocket around the BZ center (Fig. 4), which arises from the Sb orbitals. When measured at zero field, as consistent with all previous reports Liu2021; Luo2022; Kang2022; Hu2022a, the central electron pocket is nearly isotropic and circular inside the CDW state (Fig. 4a). This can be demonstrated by comparing the MDC along the vertical -K and horizontal -M directions, which shows that the separations of the two peaks are identical along the two orthogonal directions (Fig. 4b). However, when an out-of-plane magnetic field of -1.6 mT is applied, the circular Fermi pocket becomes elliptical (Fig. 4e), as can also be seen in the same comparison of the vertical (black) and horizontal (red) MDCs (Fig. 4f). Here we note that we can exclude the possibility that this Fermi surface distortion is due to potential misalignment between the sample and the magnetic coil, ellipticity of the beam, or spatial inhomogeneity of the field, as demonstrated via our extensive simulations (Supplementary Materials). Rather, the elliptical distortion of the central Fermi pocket under the application of a magnetic field suggests a field-induced RSB in CsV3Sb5 corresponding to broken symmetry. Beyond the Fermi surface, we also show the measured dispersions along the vertical and horizontal directions for both zero field and -1.6 mT (Fig. 4c,g). Their fitting shows that this elliptical distortion is not limited to the Fermi level but persists to a larger energy range of at least 250 meV below the Fermi level (Fig. 4d,h).
@@ IV Discussion
Our magneto-ARPES measurements reveal distinct field responses, inside the CDW phase, of the V 3 orbitals that make up the bands associated with the VHS of the kagome lattice and the Sb orbitals that make up the band at the BZ center. The odd-in-field response of the VHS-bands breaks C6 rotational symmetry, thus implying that the CDW phase of CsV3Sb5 is piezomagnetic, i.e., that an out-of-plane magnetic field induces an in-plane distortion. Piezomagnetism, manifesting the intrinsic coupling between TRSB and RSB, was also inferred in a recent STM study on the related compound RbV3Sb5 Xing2024, which attributed this response to a so-called congruent flux phase with coexisting CDW and loop-current order. In this state, rotational symmetry is weakly broken already in the absence of a magnetic field, and the application of a magnetic field further breaks the remaining vertical mirror, leaving intact only the horizontal mirror. Our ARPES data at zero field, however, is symmetric, and the current magneto-ARPES data cannot resolve whether a vertical mirror plane is retained or not in the presence of a field. As discussed in the Methods section and in Extended Data Fig. 6-7, this could be due to the weak character of the zero-field symmetry-breaking or more likely domain averaging within the beamspot of the ARPES measurement. As our samples are glued down, they are not free from residual strain. Considering that our beamspot is much larger than the congruent flux order domains, in the absence of field, the time-reversed pairs of loop current domains would both appear, averaging out in our ARPES signal. With the application of a field and also in the presence of residual strain, one type of time-reversed loop current domain would dominate, revealing the observed electronic response. To distinguish these scenarios from the one where symmetry is preserved at zero-field, future nano-focused synchrotron-based magneto-ARPES measurements will be helpful. Regardless of its microscopic origin, the piezomagnetic response, which implies that the CDW state spontaneously breaks TRS, is robust and strongly manifested in the V bands that give rise to the VHS, which are often associated with the emergence of unconventional correlated states in the kagome lattice Kiesel2013. A field-induced breaking of symmetry is also manifested in the bands made up of the Sb orbitals, which have been proposed to be essential for superconductivity Ritz2023_SC. Our measurements on this Fermi pocket ellipticity around the BZ center, however, do not exhibit a clear sign change upon reversing the magnetic field (Extended Data Fig. 8), in contrast to the clear odd-in-field behavior displayed by the anisotropy of the V bands.
@@
It is interesting to note that the field-response of the V 3 bands vanishes at while that of the Sb bands persists above . The origin of the latter behavior and its connection to the CDW order will require further studies. Interestingly, a recent magneto-torque measurement reported rotational symmetry breaking above that was interpreted to also break inversion symmetry Asaba2024, which is not the case in our magneto-ARPES data. One possibility is that strong fluctuations persist above and are thus “picked up” by the magnetic field, similarly to how small residual strain induces robust nematic order in the iron-based superconductors due to an enlarged nematic susceptibility Fernandes2014a. Indeed, it has been demonstrated that CsV3Sb5 is extremely sensitive to both magnetic field and strain Guo2024, and our glued sample is not free from random residual strain, which may combine with the external magnetic field to induce an ordered state. Alternatively, the magnetic field could also enhance the separation between the onsets of the CDW and loop-current order parameters in certain loop-current configurations Christensen2022; Fischer2023; Tazai2024. In either scenario, however, it is puzzling why this behavior is seen only in the Sb bands and why the anisotropy is larger outside the CDW phase than inside the CDW phase. It remains to be seen whether future experiments can map the electronic response of the distinct momentum-space regions to decoupled strain and magnetic fields, which will be helpful in disentangling different microscopic mechanisms. The new capability of magneto-ARPES demonstrated here, which enables the mapping of the spectral response in a momentum-resolved way in the presence of a tunable finite magnetic field, opens up that possibility.
V Methods
V.1 Sample growth and characterization
Single crystals of CsV3Sb5 were grown with self-flux method described in Ref. Qian2021 and Liu2024. The crystal sizes are 3 3 mm with typical residual resistivity ratio (RRR) up to 120. The resistivity measurements were performed in DynaCool (Quantum Design, Inc.) with Stanford Research SR860 Lock-in amplifiers and voltage preamplifiers.
V.2 MagnetoARPES measurements
ARPES experiments were performed using a laboratory system equipped with a helium lamp light source that is 75% linear vertical and 25% linear horizontal polarization and a Scienta DA30 electron analyzer. The angular resolution was set to 0.3∘. The total energy resolution was set to 13 meV. All the samples were cleaved in situ at around 30 K and all the measurements were conducted in an ultra-high vacuum (UHV) with a base pressure lower than 1 10-10 Torr. Results were reproduced across multiple samples.
An in situ magnetic field was applied using a magnetic coil mounted around the sample on the sample stage (Fig. 1e and f) Huang2023a. Samples were placed at the center of the magnetic coil flush with the top surface of the coil so that the magnetic field was oriented along the out-of-plane direction at the sample position. The magnitude of the field was controlled by adjusting the electric current flowing through the magnetic coil, which was characterized outside the UHV prior to the experiments Huang2023a. The magnetic field direction can be reversed by reversing the direction of the electrical current. In this experimental setup, the primary extrinsic effects of the magnetic field on ARPES measurements are constant energy contour rotation, photoelectron emission angle contraction, and momentum broadening Huang2023a. In this study, with the small applied magnetic field, only a rigid constant energy contour rotation was observed, which could be corrected post-measurements. All the ARPES measurements were performed using the DA30 deflector mode unless otherwise noted.
V.3 DFT calculations
The density functional theory-based first-principles calculations were performed with the Vienna - Simulation Package Kresse1996b. The generalized gradient approximation parameterized by Perdew, Burke, and Ernzerhof (PBE) Perdew1997 was employed for the exchange-correlation interaction between electrons throughout. The plane-wave basis set cutoff energy was 300 eV. The pristine phase of CsV3Sb5 was used to calculate the electronic structures. Notice that the crystal structure was obtained from Ref. Tan2021. Spin-orbit coupling was not involved in structural relaxation but was considered in band structure calculation. A -mesh of 12126 was utilized to sample the Brillouin zone.
VI Simulation of magneto-dichroism of ARPES spectra
We observed magneto-dichroism in the ARPES spectra between the two adjacent K regions when applying opposite magnetic fields along the -axis (Fig. 2). We have shown in the main text that the magneto-dichroism effect arises from the selective broadening of the bands at different K regions (Fig. 2, 3). The effect can be qualitatively simulated with the single-particle spectral function by manually assigning different imaginary components to the self-energy for bands around the two K regions (Extended Data Fig. 3). is described by:
is taken from the DFT calculations (dashed line in Extended Data Fig. 3). is obtained by fitting the linewidth of the corresponding experimental data. A momentum- and energy- independent is adopted for the simulation. The entire spectrum is then multiplied by a Fermi distribution at 25 K. With applying a positive magnetic field, values of 0.21 eV and 0.15 eV were individually used for , corresponding to the simulated APRES spectra around different K regions (left and right part of Extended Data Fig. 3a and d). When the magnetic field is reversed, the opposite values are used. The simulated spectra exhibit a similar magneto-dichroism effect as observed in the experimental data (Extended Data Fig. 1).
VI.1 One-step model ARPES calculations
To understand the matrix element effects in ARPES experiments under external magnetic fields, we carried out ab-initio based one-step model ARPES calculations using the spin-polarized relativistic Korringa-Kohn-Rostoker (SPRKKR) package Ebert2011Calculating under the fully relativistic four component Dirac formalism and the atomic-sphere approximation. The chosen exchange-correlation functional is based on the local spin density approximation by Vosko, Wilk, and Nusair Vosko1980Accurate. An angular momentum cutoff of l=3 was adopted for solving the KKR equations. During the self-consistent field (SCF) calculations, the Brillouin zone integration was carried out on an equidistant 29×29×14 k-mesh and the energy integral was evaluated with 32 points along the Gaussian-Legendre quadrature path. For each external magnetic field, the SCF calculations were converged separately. The Lloyd’s formula was employed to determine the Fermi level. The one-step ARPES calculations considered a semi-infinite surface model terminated by the Sb atoms with the next layer followed by the V-Sb slab, rather than Cs atoms. The surface barrier takes the z-dependent Rundgren-Malmström type Rundgren1977Transmission. Experimental light incidence angles and the photon energy of 21.2 eV were adopted in the calculations. To simulate the mixed polarization of the helium lamp, the final calculated one-step ARPES spectra were obtained through a linear superposition of 25% p-polarization and 75% s-polarization results. The life-time effects of the initial and final states were simulated by an imaginary potential of 0.05 eV and 1.5 eV, respectively. The external magnetic field theoretically enters all three parts of the photoemission matrix elements—initial states, the relativistic dipole operator, and the spin-polarized time-reversed low-energy electron diffraction final states, the treatment of which has been described previously for ferromagnetic materials Braun1996Theory. The classical photoelectron deflection in the vacuum due to the Lorentz force under magnetic fields was not considered in the one-step model ARPES calculations.
The results of the calculations are summarized in Extended Data Fig. 5. First, we show the zero-field constant energy contour (CEC) 260 meV below the Fermi level (Extended Data Fig. 5a). Notice that the Fermi level in the calculations is approximately 90 meV higher than the one in experiments. Therefore, the theoretical CEC in Extended Data Fig. 5a presents reasonable resemblance to the experimental – = -0.35 eV CEC of Fig. 2(a) taken slightly below the Dirac cone. Without adjusting the theoretical Fermi level, we extract the momentum slice Cut 1 at Å-1 and compare its band dispersions and MDCs under opposite directions of large external fields around 117.5 T. The results are summarized in Extended Data Fig. 5b-d. The horizontal dashed lines near the MDC (taken at – = -0.2 eV) inner peaks clearly indicate a left-right symmetric feature at 0 T, but field-odd asymmetric behavior under the applied large magnetic fields. While this may seem consistent with the experimental data in Fig. 2e-g, the fields required to observe this asymmetry in the simulations are five orders of magnitude larger than the experimental ones. The difference in the theoretical calculations here is that the TRS-breaking comes from the large magnetic fields and the induced moments on V atoms, treated within the DFT. To gain further insights on the asymmetric MDC peaks in Extended Data Fig. 5c-d, we present the spin-resolved MDCs under different magnetic fields in Extended Data Fig. 5e-g with the projection axis chosen along z. Since if only considering the out-of-plane magnetic field which breaks the time reversal symmetry and all vertical mirrors, the magnetic mirror (-K containing) preserves, protecting the band degeneracy between and . Therefore, we infer that spin-dependent multiple scattering and relativistic effects could contribute to the asymmetric MDC line-shapes with respect to the -M plane. However, this only becomes discernible in the calculations with magnetic fields nearly five orders of magnitude larger than the experimental ones. As a comparison, we show in Extended Data Fig. 5h the Cut 1 spin-integrated MDCs taken at – = -0.2 eV under 0, +2.35, and -2.35 T fields. The calculated results clearly overlap with each other without any signatures of the magneto-dichroism observed in ARPES experiments.
VI.2 Model for the fingerprints of piezomagnetism in ARPES
We discuss a phenomenological model to shed light on the manifestations of piezomagnetism in our magneto-ARPES data. Specifically, we consider the congruent CDW flux phase proposed in Ref. Xing2024, which has the non-zero piezomagnetic tensor component , defined as , where is the strain tensor and is the magnetic field. Thus, implies that a non-zero shear strain is induced by an out-of-plane field . Importantly, the congruent CDW flux phase also displays a non-zero distortion in the absence of a magnetic field. Our goal is to model the impact of both effects on the V and Sb Fermi surfaces.
We start with the large vanadium kagome Fermi surface. For simplicity, we model it in terms of a simple tight-binding model on the kagome lattice. Denoting the three kagome sublattices as A, B, and C, and following the coordinate system of Ref. Guo2009a, the nearest-neighbor tight-binding model is given by:
where and:
Here, denotes the chemical potential, is the hopping parameter and , , . The band structure has a saddle point at = 0, resulting in a perfectly hexagonal Fermi surface.
In the congruent CDW flux phase proposed in Ref. Xing2024, two additional terms emerge:
with:
Here, and are parameters that transform as the and form factors, respectively, and is the out-of-plane magnetic field. Note that this parametrization gives one of the domains only. While both terms individually break symmetry, their combination also breaks all in-plane twofold rotation axes (i.e., all vertical mirrors). We can diagonalize to obtain the distorted Fermi surface under a magnetic field. We note that because the congruent CDW flux phase breaks translational symmetry, it also folds the band structure. We do not include these effects since our goal is to focus on the piezomagnetic effect.
The fact that the ARPES data without a magnetic field is symmetric indicates that either or , as shown in the Extended Data Fig. 6. For the calculations incorporating a non-zero magnetic field, we adopted .
To visualize the behavior of the Fermi surface in the presence of a magnetic field with this model, we set (in order to be slightly away from perfect nesting) and calculate the Fermi surface with the parameters (, , ) = (1, 0.01, 0.1) to represent a positive magnetic field and (, , ) = (1, 0.01, -0.1) to represent a negative magnetic field (Extended Data Fig. 7). The Fermi surfaces obtained in the presence of a magnetic field for both situations preserve inversion symmetry but break the mirror symmetry with respect to the horizontal –M plane (Extended Data Fig. 7b and c). This is consistent with the symmetry-breaking behavior observed in the experimental data, which however is manifested in the spectral weight distribution of the Fermi surface.
For the case of the small Sb Fermi pocket, it suffices to model the electronic dispersion in terms of a parabolic dispersion with two broken-symmetry terms analogous to and above in Eq. (17):
where is the angle with respect to the axis, which corresponds to a direction. We use this equation to fit the angle dependence of the Fermi surface diameter extracted from the ARPES data. To simplify the notation, we rewrite the equation above as follows:
To accurately capture the Fermi surface distortion under a magnetic field, we first extract the Fermi surface diameter as a function of the Fermi surface angle by fitting the Fermi surface mappings obtained with and without a magnetic field (Extended Data Fig. 8a). We then normalize the Fermi surface diameter under a magnetic field by dividing it by the diameter measured without a magnetic field, thus excluding potential extrinsic effects. The normalized Fermi surface diameter as a function of the Fermi surface angle exhibits a more pronounced two-fold rotational symmetry (Fig. 4j). We then fit the Fermi momentum using Eq. 19 (Extended Data Fig. 8b), yielding fitting parameters and . The fitting results reveal a consistent angular deviation of the longitudinal axis of the elliptical Fermi surface from the BZ high-symmetry direction, in agreement with the fitting approach used in the main text.
VII Data Availability
All data needed to evaluate the conclusions are present in the paper and extended data. Additional data are available from the corresponding authors upon reasonable request.
VIII Code Availability
The band structure calculations used in this study are available from the corresponding authors upon reasonable request.
IX Acknowledgments
The ARPES work at Rice University was supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant No. GBMF9470 and the Robert A. Welch Foundation Grant No. C-2175. The single-crystal synthesis work at UW was supported by the Air Force Office of Scientific Research (AFOSR) under Award No. FA2386-21-1-4060 and the David Lucile Packard Foundation. R.M.F. was supported by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0423. Z.W. is supported by the Department of Energy, Basic Energy Sciences Grant No. DE-FG02-99ER45747.
X Author contributions
M.Y. and J.Huang proposed and designed the project. Z.L. grew the CsV3Sb5 single crystals with the help of J.M.D. under the guidance of J.H.C.. J.Huang, Z.R., J.Hyun, T.H. and Z.Y. carried out the ARPES measurements under the guidance of M.Y. Y.H. and J.K.. J.Huang performed the magneto-ARPES data analysis. Y.X., Z.L., and J.M.D. grew single crystals for control measurements under the guidance of P.D. and J.H.C.. H.T., B.Y. and Z.W. conducted the first-principles calculations. Y.Z. carried out the ab-initio based one-step model ARPES calculations with advice from A.P. and J.M.. Y.Z. analyzed the one-step calculated data. R.M.F. performed the phenomenological theoretical analysis. J.Huang and M.Y. wrote the paper with input from all co-authors.
XI Competing interests
The authors declare that they have no competing interests.

### v2

@@ III Field response of the Sb 5 band at
Next, we turn our attention to the electron pocket around the BZ center (Fig. 4), which arises from the Sb orbitals. When measured at zero field, as consistent with all previous reports Liu et al. (2021); Luo et al. (2022); Kang et al. (2022); Hu et al. (2022), the central electron pocket is nearly isotropic and circular inside the CDW state (Fig. 4a). This can be demonstrated by comparing the MDC along the vertical -K and horizontal -M directions, which shows that the separations of the two peaks are identical along the two orthogonal directions (Fig. 4b). However, when an out-of-plane magnetic field of -1.6 mT is applied, the circular Fermi pocket becomes elliptical (Fig. 4e), as can also be seen in the same comparison of the vertical (black) and horizontal (red) MDCs (Fig. 4f). Here we note that we can exclude the possibility that this Fermi surface distortion is due to potential misalignment between the sample and the magnetic coil, ellipticity of the beam, or spatial inhomogeneity of the field, as demonstrated via our extensive simulations (Supplementary Materials). Rather, the elliptical distortion of the central Fermi pocket under the application of a magnetic field suggests a field-induced RSB in CsV3Sb5 corresponding to broken symmetry. Beyond the Fermi surface, we also show the measured dispersions along the vertical and horizontal directions for both zero field and -1.6 mT (Fig. 4c,g). Their fitting shows that this elliptical distortion is not limited to the Fermi level but persists to a larger energy range of at least 250 meV below the Fermi level (Fig. 4d,h).
@@ IV Discussion
Our magneto-ARPES measurements reveal distinct field responses, inside the CDW phase, of the V 3 orbitals that make up the bands associated with the VHS of the kagome lattice and the Sb orbitals that make up the band at the BZ center. The odd-in-field response of the VHS-bands breaks C6 rotational symmetry, thus implying that the CDW phase of CsV3Sb5 is piezomagnetic, i.e., that an out-of-plane magnetic field induces an in-plane distortion. Piezomagnetism, manifesting the intrinsic coupling between TRSB and RSB, was also inferred in a recent STM study on the related compound RbV3Sb5 Xing et al. (2024), which attributed this response to a so-called congruent flux phase with coexisting CDW and loop-current order. In this state, rotational symmetry is weakly broken already in the absence of a magnetic field, and the application of a magnetic field further breaks the remaining vertical mirror, leaving intact only the horizontal mirror. Our ARPES data at zero field, however, is symmetric, and the current magneto-ARPES data cannot resolve whether a vertical mirror plane is retained or not in the presence of a field. As discussed in the Methods section and in Extended Data Fig. 6-7, this could be due to the weak character of the zero-field symmetry-breaking or more likely domain averaging within the beamspot of the ARPES measurement. As our samples are glued down, they are not free from residual strain. Considering that our beamspot is much larger than the congruent flux order domains, in the absence of field, the time-reversed pairs of loop current domains would both appear, averaging out in our ARPES signal. With the application of a field and also in the presence of residual strain, one type of time-reversed loop current domain would dominate, revealing the observed electronic response. To distinguish these scenarios from the one where symmetry is preserved at zero-field, future nano-focused synchrotron-based magneto-ARPES measurements will be helpful. Regardless of its microscopic origin, the piezomagnetic response, which implies that the CDW state spontaneously breaks TRS, is robust and strongly manifested in the V bands that give rise to the VHS, which are often associated with the emergence of unconventional correlated states in the kagome lattice Kiesel et al. (2013). A field-induced breaking of symmetry is also manifested in the bands made up of the Sb orbitals, which have been proposed to be essential for superconductivity Ritz et al. (2023). Our measurements on this Fermi pocket ellipticity around the BZ center, however, do not exhibit a clear sign change upon reversing the magnetic field (Extended Data Fig. 8), in contrast to the clear odd-in-field behavior displayed by the anisotropy of the V bands.
@@
It is interesting to note that the field-response of the V 3 bands vanishes at while that of the Sb bands persists above . The origin of the latter behavior and its connection to the CDW order will require further studies. Interestingly, a recent magneto-torque measurement reported rotational symmetry breaking above that was interpreted to also break inversion symmetry Asaba et al. (2024), which is not the case in our magneto-ARPES data. One possibility is that strong fluctuations persist above and are thus “picked up” by the magnetic field, similarly to how small residual strain induces robust nematic order in the iron-based superconductors due to an enlarged nematic susceptibility Fernandes et al. (2014). Indeed, it has been demonstrated that CsV3Sb5 is extremely sensitive to both magnetic field and strain Guo et al. (2024a), and our glued sample is not free from random residual strain, which may combine with the external magnetic field to induce an ordered state. Alternatively, the magnetic field could also enhance the separation between the onsets of the CDW and loop-current order parameters in certain loop-current configurations Christensen et al. (2022); Wagner et al. (2023); Tazai et al. (2024). In either scenario, however, it is puzzling why this behavior is seen only in the Sb bands and why the anisotropy is larger outside the CDW phase than inside the CDW phase. It remains to be seen whether future experiments can map the electronic response of the distinct momentum-space regions to decoupled strain and magnetic fields, which will be helpful in disentangling different microscopic mechanisms. The new capability of magneto-ARPES demonstrated here, which enables the mapping of the spectral response in a momentum-resolved way in the presence of a tunable finite magnetic field, opens up that possibility.